equired information [The following information applies to the questions displayed below.] Tyrell Co. entered into the following transactions involving short-term liabilities in 2017 and 2018. 2017 Apr. 20 Purchased $35,500 of merchandise on credit from Locust, terms n/30. Tyrell uses the perpetual inventory system. May 19 Replaced the April 20 account payable to Locust with a 90-day, $35,000 note bearing 8% annual interest along with paying $500 in cash. July 8 Borrowed $63,000 cash from NBR Bank by signing a 120-day, 11% interest-bearing note with a face value of $63,000. __?__ Paid the amount due on the note to Locust at the maturity date. __?__ Paid the amount due on the note to NBR Bank at the maturity date. Nov. 28 Borrowed $33,000 cash from Fargo Bank by signing a 60-day, 9% interest-bearing note with a face value of $33,000. Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank. 2018 __?__ Paid the amount due on the note to Fargo Bank at the maturity date.

2. Determine the interest due at maturity for each of the three notes. (Do not round your intermediate calculations. Use 360 days a year.)
Answer #1
2
Principal X Rate X Time = Interest
Locust 35000 X 8% X 90/360 = 700
NBR Bank 63000 X 11% X 120/360 = 2310
Fargo Bank 33000 X 9% X 60/360 = 495